OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2).

OpenStudy (anonymous):

\[f(\Box)=-4\times \Box -1\] put \(-2\) in the box

OpenStudy (anonymous):

= 7

OpenStudy (anonymous):

\[f(-2)=-4\times (-2)=8-1=7\]

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

replace \(x\) by \(-2\) your first step it to write \[-4\times (-2)-1\]

OpenStudy (anonymous):

ohhh i got it now

OpenStudy (anonymous):

then \(-4\times (-2)=8\) so you get \(8-1\) and then \(8-1=7\)

OpenStudy (anonymous):

it is very easy to solve just substitute..